# Self-Pro: Self-Prompt and Tuning Framework for Graph Neural Networks

### v2

@@ 5 Experiments
$\bullet$(RQ1) How effective is Self-Pro for node classification on homophilic graphs and heterophilious graphs under the few-shot setting?
$\bullet$(RQ2) How adaptable is Self-Pro when transferred to different tasks and shots?
@@ 5.1 Experimental Settings
We conduct experiments on homophilic and heterophilic graphs. For homophilic graphs, we adopt citation networks: Cora, Citeseer and Pubmed [12], and co-purchase graphs: Computer and Photo [12]. For heterophilic graphs, we adopt webpage networks: Texas, Wisconsin, Cornell [22], and Wikipedia page–page networks: Chameleon, Squirrel and Crocodile [23]. Further details of the datasets are summarized in Appendix 0.B.
5.1.2 Setup and evaluation.
For node classification, we follow the $k$-shot setup in [20]. Specifically, we randomly sample 1 node per class for training, 5 nodes per class for validation and the remaining nodes for testing. We run the experiments 10 times, and report mean accuracy with standard deviation. For link prediction, we follow the setup in [15] and construct the validation/test set by randomly selecting 20%/10% edges from the original graph. After removing selected edges, we train the model with the remaining 70% edges. We run the experiments 10 times and use AUC and AP to report the average results with standard deviation.
5.1.3 Baselines.
To evaluate Self-Pro, we compare it with the state-of-the-art methods from three categories. More baseline details can be found in Appendix C.
$\bullet$ Graph prompt methods: GPPT [24], GraphPrompt [20] and ProG [25].
@@ 5.2 Few-shot Node Classification (RQ1)
For node classification, we inject either structural or semantic prompt separately and report the best results. More experiments and analysis about different prompts are described in Section 5.4 and Appendix 0.C. Table 1 illustrates the results of few-shot on homophilic graphs. These observations are as following:
(1) Supervised GNN models achieve poor performance in most cases, demonstrating heavy dependency on labeled data. However, some pre-trained models achieve even worse performance than supervised models on Pubmed, indicating the negative transfer of pre-training knowledge.
(2) Self-Pro performs the best across all datasets and other prompt methods such as ProG also achieve good results in few-shot setting, demonstrating that unified task template can fully leverage the capabilities of pre-trained models.
We also present the results of few-shot node classification on heterophilic graphs in Table 2:
(3) Supervised methods such as H2GCN, GloGNN and self-supervised methods such as MUSE and GraphACL designed for heterophily achieve good results due to the ability to capture higher-order similarities beyond local neighborhood. However, GAE, GraphPrompt, DGI and GRACE achieve poor performance on heterophilic datasets since they follow the homophily assumption.
(4) MLP achieves good results on Cornell, Texas and Wisconsin, indicating that the semantic of node attributes plays a vital role in heterophilic graph.
@@
Table 3: Link prediction results. The best and second best performance for each dataset are marked with bold and underline.
@@ 5.3 Performance under different tasks and shots (RQ2)
5.3.2 Performance under different shots.
We study the impact of shot number on the Pubmed and Chameleon in Fig. 3. We vary the number of shots between 1 and 10, and compare with competitive baselines. Note that 10 shots per class might be sufficient for semi-supervised node classification. We can see that Self-Pro consistently outperforms the baselines on Pubmed and Chameleon. The performance under the semi-supervised setting can be seen in Appendix 0.D.
(a) Pubmed
(b) Chameleon
@@ 5.4 Ablation Study (RQ3)
We still follow the above setting of few-shot node classification task, and more details and analysis can be found in Appendix 0.C. In Fig.4 we can see the results:
(1) The degradation without prompt template mainly stems from the negative transfer problem, especially when labeled nodes are limited.
(2) The decline without tuning indicates that the projector plays a crucial role in the knowledge transfer between pre-training and downstream tasks.
(3) Prompt injection improves the performance, indicating that providing extra information based on graph itself is beneficial for tasks with limited labels and the projector can effectively integrate prompt that does not appear during the pre-training phase to enhance the performance in downstream tasks.
@@ 5.5 Parameter Analysis (RQ4)
We end this section with a sensitivity analysis on hyper-parameters in Self-Pro, i.e., the weight of prompt injection $\mu$. Meanwhile, we demonstrate the results using self weight $w$ mentioned in Section 4.4. Specificlly, we conduct experiments on node classification by varying $\mu$ from 0 to 1 to compare with the self weight strategy. Fig. 5 illustrates the accuracy under different $\mu$ on Cora, Citeseer and Texas. We see that self weight strategy and smaller injection weights can assist in node classification on homophilic graph, while larger injection ratios are beneficial for heterophilic graphs where the semantic information is more crucial. More experimental results can be seen in Appendix 0.E.
@@ 6 Conclusion
In this paper, we proposed Self-Pro, a framework utilizing self-prompting and tuning mechanism for graphs. The framework incorporated asymmetric graph contrastive learning to capture complex patterns for graphs with homophily and heterophily and align the pretext and downstream tasks into the same template. Then we reused the projector in the pre-training phase as the downstream adapter for efficient tuning and injected prompts from semantic and structure perspectives based on the information inherented in the graph itself. Finally, our experimental results demonstrated the effectiveness of Self-Pro.
Appendix 0.A Model Analysis
0.A.1 Time Complexity
We analyze the time complexity of main components in our model. Assume an input graph has $n$ nodes and $m$ edges, let the GNN encoder contain $l$ layers. During the pre-training phrase, the time complexities of GNN encoder is $O(dml+nd^{\prime}dl)$, where $d$ is the dimensionality of initial features, $d^{\prime}$ is the dimensionality of final representations, and the time complexities of projector is $O(nd^{\prime 2})$ since we implement it using a simple linear layer. Let $b$ be the number of selected samples in a batch, the complexity of contrastive loss is $O(nbd^{\prime})$. During prompt tuning, the time complexity for $k$-shot classification task with $c$ classes is $O(kc^{2}d^{\prime})$, and the time complexity for link prediction is $O(nd^{\prime})$. During the inference phrase, the time complexity for classification task is $O(cd^{\prime})$, and the time complexity for link prediction is $O(d^{\prime})$.
0.A.2 Parameter Efficiency
We also compare the number of parameters to be introduced for node classification task with a few representative models in Table 4. GRACE and MVGRL employ a linear classifier for downstream tuning, while GraphPrompt introduces a prompt vector which shares the same dimension with the node representation for prompt tuning. For proposed Self-Pro, it outperforms the baselines GRACE, MVGRL and GraphPrompt without additional parameters as we have seen earlier, demonstrating its superiority.
Table 4: Study of parameter efficiency on node classification.
Method Cora Citeseer Pubmed Computer Photo
MVGRL 3,591 3,078 1,539 4,104 5,130
GRACE 1,799 1,542 1,028 2,056 2,570
GraphPrompt 256 256 256 256 256
Self-Pro 0 0 0 0 0
Appendix 0.B Further Details of Datasets
0.B.1 Dataset Statistics
Statistics of datasets in the experiments are summarized in Table 5.
Table 5: Dataset statistics.
Dataset Classes Nodes Edges Features $\mathbf{h}(\mathcal{G})$ $\mathbf{h}(\mathcal{G}_{2})$
Cora 7 2,708 5,429 1,433 0.81 0.71
Citeseer 6 3,327 4,732 3,703 0.74 0.56
Pubmed 3 19,717 44,338 500 0.80 0.74
Photo 8 7,650 119,081 745 0.83 0.66
Computer 10 13,752 574,418 767 0.78 0.55
Texas 5 183 309 1,703 0.11 0.54
Cornell 5 183 295 1,703 0.30 0.40
Wisconsin 5 251 499 1,703 0.21 0.42
Chameleon 5 2,277 36,101 2,325 0.23 0.35
Squirrel 5 5,201 216,933 2,089 0.22 0.22
Crocodile 5 11,631 360,040 2,089 0.25 0.30
One-hop Homophily ratio.
We use the edge homophily [22] to measure the one-hop neighbor homophily of the graph. The one-hop homophily ratio $\mathbf{h}(\mathcal{G})$ is defined as the fraction of edges connecting nodes with the same label:
$\mathbf{h}(\mathcal{G})=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in\mathcal{V}}\frac{|\{u\in\mathcal{N}(v):y_{v}=y_{u}\}|}{|\mathcal{N}(v)|}.$ (15)
Two-hop Homophily ratio.
We use the monophily level $\mathbf{h}(\mathcal{G}_{2})$ [34] to measure the two-hop similarities (monophily) of the graph. The two-hop homophily ratio is defined as:
$\mathbf{h}(\mathcal{G}_{2})=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in\mathcal{V}}\frac{|\{u\in\mathcal{N}_{2}(v):y_{v}=y_{u}\}|}{|\mathcal{N}_{2}(v)|},$ (16)
where $\mathcal{N}_{2}(v)$ denotes the set of two-hop neighbors of node $v$.
0.B.2 Dataset Details
Cora, Citeseer, and Pubmed.
They are four citation networks, where nodes represent publications and edges are citation links. Features for each node are the keywords it contains. Each dimension in the feature vector indicates the presence of a keyword in the publication. Nodes in these datasets are associated with labels that describe research topics of publications.
Computer and Photo.
They are two co-purchase graphs obtained from Amazon, where nodes represent products and and edges connect frequently bought products. The node features represent product reviews, while the class labels indicate the product category.
Texas, Wisconsin and Cornell.
They are webpage networks collected by Carnegie Mellon University from computer science departments of different universities. In each network, nodes represent web pages, and edges indicate hyperlinks between web pages. The node features consist of bag-of-words representations of the web pages. The objective is to classify the nodes into five categories: student, project, course, staff, and faculty.
Chameleon, Crocodile and Squirrel.
These networks are derived from Wikipedia, where nodes represent web pages and edges represent hyperlinks connecting them. The node features consist of several informative nouns extracted from the Wikipedia pages. The node labels are assigned based on the average daily traffic of the corresponding web pages.
Appendix 0.C Analysis about Prompts
In Fig. 6, we present more experimental details about the prompt design and different prompts in Self-Pro. Consistent with our previous conclusion, incorporating two-order neighbors as prompts improves performance on homopilous graphs such Citeseer and Pubmed. Upon examining Table 5, we observed that on these two datasets, a majority of the nodes have the same labels as their two-order neighbors. This also explains why we can freeze the parameters of the pre-trained model and directly modify the graph structure to incorporate two-hop neighbor information, thereby optimizing the performance of node classification. However, if structural prompts are applied to graphs where the one-hop homophily or even two-hop homophily is low, it can lead to a decrease in performance, such as on Squirrel. This is because a chaotic structure cannot provide effective cues to the model. Meanwhile, we find that semantic prompts can improve performance across various datasets. This is because when the downstream task has a small number of labels, providing semantic information about node features can provide more references for the node classification task.
(a) Citeseer
(b) Pubmed
(c) Squirrel
(d) Wiscousin
(e) Texas
(f) Chameleon
Figure 6: More details about the impact of different prompts.
Appendix 0.D Performance under Semi-Supervised Setting
We further investigate the performance of Self-Pro under the semi-supervised setting with scarce-labeled training data. For homophilic graphs, we adopt the public splits with 20 nodes per class for training, 500 nodes for validation and 1,000 nodes for testing [12, 35]. For heterophilic graphs, we adopt the commonly used training/validation/test split ratio of 48/32/20 as previous works [22, 35]. The results are shown in Table 6 and we can find that:
(1) When labels are sufficient, GraphACL and MUSE demonstrate impressive performance on both homophilic and heterophilic graphs. This can be attributed to their frameworks that adequately consider the heterophily of graphs.
(2) Self-Pro still outperforms traditional supervised models (like GCN, GAT and MLP) and self-supervised GNN models (like DGI, GRACE, MVGRL, MUSE and GraphACL) across most datesets, indicating that our framework can effectively transfer pre-trained knowledge to downstream tasks once again.
(3) Most baselines perform even worse than simple MLP on heterophilic graphs, indicating the importance of feature semantics. This also confirms that our model can handle the heterophily on graphs through semantic prompt injection.
[…]
Method Cora Citeseer Pubmed Texas Cornell
MLP 56.11±0.34 56.91±0.46 71.35±0.05 81.62±3.51 81.08±0.98
GCN 81.58±0.12 70.30±0.21 78.97±0.35 60.00±2.85 57.03±2.30
GAT 83.03±0.24 72.37±0.42 79.03±0.32 61.62±3.87 78.63±5.92
DGI 82.37±0.62 71.94±0.78 79.43±0.84 60.59±2.56 63.35±1.61
GRACE 83.01±0.53 72.41±0.38 81.15±0.34 58.57±2.68 55.86±1.95
MVGRL 83.03±0.37 72.84±0.64 79.95±0.38 63.83±5.16 64.08±5.73
MUSE 82.24±0.47 71.14±0.82 81.90±0.59 83.73±1.39 82.09±2.58
GraphACL 84.20±0.31 73.58±0.41 82.20±0.73 71.08±0.57 59.95±2.23
Self-Pro 84.58±0.18 73.15±0.68 82.11±0.59 85.14±0.63 86.29±0.87
Appendix 0.E Analysis about Hyper-parameter
In this section, we present more details in Fig. 7 about the impact of hyper-parameters including the weight of prompt injection $\mu$ to compare with the self weight strategy.
(a) Pubmed
(b) Wisconsin
(c) Cornell
(d) Crocodile
Figure 7: Sensitivity analysis on the injection weight of prompts.
We find that a smaller proportion of semantic prompt injection can help with node classification on the Pubmed and Crocodie, while a larger proportion of semantic prompt injection can improve performance on the Wisconsin and Cornell.
Appendix 0.F Further Details of Baselines
In this section, we present more details for the baselines, which are chosen from three main categories:
Supervised graph neural networks.
$\bullet$ GCN [12]: GCN resorts to mean-pooling based neighborhood aggregation to receive messages from the neighboring nodes for node representation learning in an end-to-end manner.
$\bullet$ GAT [28]: GAT also depends on neighborhood aggregation for node representation learning in an end-to-end manner, while it can assign different attention to neighbors to reweigh their weights.
$\bullet$ H2GCN [39]: H2GCN employs the ego- and neighbor-embedding separation to combine the representations of higher-order neighborhoods, which is specially designed for graphs with heterophily.
$\bullet$ GloGNN [39]: GloGNN proposes to go beyond local homophily and discover additional connections from a global perspective to mitigate heterophily. It treats the input graph into a fully-connected one and efficiently generates a node’s embedding from the global view.
Graph pre-training models.
$\bullet$ GAE [13]: GAE encodes the input node features into the latent space and then reconstructs the adjacent matrix to obtain the representations of nodes. By minimizing the reconstruction loss, GAE(VGAE) learns compact representations that capture important graph properties.
$\bullet$ MaskGAE [14]: MaskGAE adopts Masked Graph Modeling (MGM) as a principled pretext: masking a portion of edges and attempting to reconstruct the missing part with partially visible, unmasked graph structure.
$\bullet$ SeeGera [15]: SeeGera adopts a hierarchical variational framework, and mainly focuses on feature reconstruction and structure/feature masking. On the one hand, SeeGera co-embeds both nodes and features in the encoder and reconstructs both links and features in the decoder. On the other hand, SeeGera adds an additional layer for structure/feature masking, which boosts the model generalizability.
$\bullet$ DGI [29]: DGI capitalizes on a self-supervised method for pre-training, which is based on the concept of mutual information. It maximizes the mutual information between the local augmented instances and the global representation.
$\bullet$ GRACE [40]: GRACE generates two graph views by corruption and learn node representations by maximizing the agreement of node representations in these two views.
$\bullet$ MVGRL [6]: MVGRL learning node and graph level representations by contrasting encodings from two structural views of graphs including first-order neighbors and a graph diffusion.
$\bullet$ MUSE [35]: MUSE is a graph constrastive learning framework for both homophilic and heterophilic graphs. It captures the information from semantic and contextual views and fuse the node representations with an information fusion controller.
$\bullet$ GraphACL [34]: The key idea of GraphACL is to capture both a local neighborhood context of one-hop and a monophily similarity of two hops in one single objective. Specifically, we present a simple asymmetric contrastive objective for node representation learning.
Graph prompt models.
$\bullet$ GPPT [24]: GPPT pre-trains a GNN model based on the link prediction task, and employs a learnable prompt to reformulate the downstream node classification task into the same format as link prediction.
$\bullet$ GraphPrompt [20]: GraphPrompt introduces a unification framework by mapping different tasks to subgraph similiar computation and proposed a learnable task-specific prompt vector that guides each downstream task to make full of the pre-trained model.
$\bullet$ ProG [25]: ProG propose a multi-task prompting method for graph models. They first unify the format of graph prompts and language prompts with the prompt token, token structure, and inserting pattern. Then, to further narrow the gap between various graph tasks, they further reformulate downstream problems to the graph-level. They also introduce meta-learning for more reliable and general for different tasks.
Appendix 0.G Further Implementation Details
We implement our model by PyTorch and optimize the model by Adam optimizer and run experiments on a machine with a NVIDIA RTX A800 GPU. In the pre-training phase, for our backbone GraphACL, we adjust the learning rate within {0.0001, 0.0005, 0.001, 0.005, 0.01}, the dimension of node representation $D$ within {128, 256, 512, 1024, 2048, 4096, 8192}, the temperature $\tau$ within {0.25, 0.5, 0.75, 0.99, 1} and the weight decay within {0, 0.0001, 0.0003, 0.000001}. In the prompt tuning phase, we tune the learning rate within {0.001, 0.005, 0.01, 0.05} and the weight of prompt injection $\mu$ within {0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9} if used for node classification. For link prediction, we use dot-product operator in SeeGera [15] as the decoder to reconstruct the graph structure and we do not inject any semantic or structural prompts to demonstrate the powerful capability of prompt tuning.

### v3

@@ 5 Experiments
$\bullet$(RQ1) How effective and efficient is Self-Pro for node classification on homophilous and heterophilous graphs under the few-shot setting?
$\bullet$(RQ2) How adaptable is Self-Pro when transferred to different shots and tasks?
@@ 5.1 Experimental Settings
We conduct experiments on homophilous and heterophilous graphs. For graphs with homophily, we adopt citation networks: Cora, Citeseer and Pubmed [12], and co-purchase graphs: Computer and Photo [12]. For graphs with heterophily, we adopt webpage networks: Texas, Wisconsin, Cornell [22], and Wikipedia page–page networks: Chameleon, Squirrel and Crocodile [23]. Further details of the datasets are summarized in Table 1.
Table 1: Statistics of datasets used in experiments.
Dataset Classes Nodes Edges Features
Cora 7 2,708 5,429 1,433
Citeseer 6 3,327 4,732 3,703
Pubmed 3 19,717 44,338 500
Photo 8 7,650 119,081 745
Computer 10 13,752 574,418 767
Texas 5 183 309 1,703
Cornell 5 183 295 1,703
Wisconsin 5 251 499 1,703
Chameleon 5 2,277 36,101 2,325
Squirrel 5 5,201 216,933 2,089
Crocodile 5 11,631 360,040 2,089
@@
5.1.2 Setup and evaluation.
For node classification, we follow the $k$-shot setup in [20]. Specifically, we randomly sample 1 node per class for training, 5 nodes per class for validation and the remaining nodes for testing. We run the experiments 10 times, and report mean accuracy with standard deviation. For link prediction, we follow the setup in [15] and construct the validation/test set by randomly selecting 20%/10% edges from the original graph. After removing the selected edges, we train the model with the remaining 70% edges and use the dot-product decoder. We run the experiments 10 times and use two commonly used metrics, the area under the ROC curve (AUC) and the average precision (AP), to report the average results with standard deviation.
5.1.3 Baselines.
To evaluate Self-Pro, we compare it with the state-of-the-art methods from three categories:
$\bullet$ Graph prompt methods: GPPT [24], GraphPrompt [20] and ProG [25].
@@ 5.2 Few-shot Node Classification (RQ1)
For node classification, we inject either structural or semantic prompt separately and report the best results. More experiments and analysis about different prompts will be elaborated on in Section 5.4.
5.2.1 Few-shot on homophilous graphs.
Table 2 illustrates the results of few-shot node classification on homophilous graphs. These observations are as following:
(1) Supervised GNN models achieve poor performance in most cases, demonstrating heavy dependency on labeled data. However, some pre-trained models achieve even worse performance than supervised models on Pubmed, indicating the negative transfer of pre-trained knowledge.
(2) Self-Pro performs the best across all the datasets and other prompt methods such as ProG also achieve good results in few-shot setting, demonstrating that a unified task template can fully leverage the capabilities of pre-trained models.
5.2.2 Few-shot on heterophilous graphs.
We also present the results of few-shot node classification on heterophilous graphs in Table 3. These observations are as following:
(3) Supervised methods such as H2GCN, GloGNN and self-supervised methods such as MUSE and GraphACL designed for heterophily achieve good results due to the ability to capture higher-order similarities beyond local neighborhood. However, GAE, GraphPrompt, DGI and GRACE achieve poor performance on heterophilous datasets since they follow the homophily assumption.
(4) MLP achieves good results on Cornell, Texas and Wisconsin, indicating that the semantic of node attributes plays a vital role on heterophilous graph.
@@
Table 4: Study of parameter efficiency on node classification.
Method Cora Citeseer Pubmed Computer Photo
MVGRL 3,591 3,078 1,539 4,104 5,130
GRACE 1,799 1,542 1,028 2,056 2,570
GraphPrompt 256 256 256 256 256
Self-Pro 0 0 0 0 0
5.2.3 Parameter efficiency.
We also compare the number of parameters to be introduced in Self-Pro for node classification task with a few representative models in Table 4. GRACE and MVGRL employ a linear classifier for downstream tuning, while GraphPrompt introduces a prompt vector which shares the same dimension with the node representation for prompt tuning. For proposed Self-Pro, it outperforms the baselines GRACE, MVGRL and GraphPrompt without additional parameters as we have seen earlier, demonstrating its superiority.
(a) Pubmed
(b) Chameleon
[…]
Method Cora Citeseer Pubmed Texas Cornell
MLP 56.11±0.34 56.91±0.46 71.35±0.05 81.62±3.51 81.08±0.98
GCN 81.58±0.12 70.30±0.21 78.97±0.35 60.00±2.85 57.03±2.30
GAT 83.03±0.24 72.37±0.42 79.03±0.32 61.62±3.87 78.63±5.92
DGI 82.37±0.62 71.94±0.78 79.43±0.84 60.59±2.56 63.35±1.61
GRACE 83.01±0.53 72.41±0.38 81.15±0.34 58.57±2.68 55.86±1.95
MVGRL 83.03±0.37 72.84±0.64 79.95±0.38 63.83±5.16 64.08±5.73
MUSE 82.24±0.47 71.14±0.82 81.90±0.59 83.73±1.39 82.09±2.58
GraphACL 84.20±0.31 73.58±0.41 82.20±0.73 71.08±0.57 59.95±2.23
Self-Pro 84.58±0.18 73.15±0.68 82.11±0.59 85.14±0.63 86.29±0.87
5.3 Performance under different shots and tasks (RQ2)
5.3.1 Performance under different shots.
We study the impact of shot number on the Pubmed and Chameleon in Fig. 3. We vary the number of shots between 1 and 10, and compare with competitive baselines. Note that 10 shots per class might be sufficient for semi-supervised node classification. We can see that Self-Pro consistently outperforms the baselines on Pubmed and Chameleon. Further, we investigate the performance of Self-Pro under the semi-supervised setting. For homophilous graphs, we adopt the public splits with 20 nodes per class for training, 500 nodes for validation and 1,000 nodes for testing [12, 35]. For heterophilous graphs, we adopt the commonly used training/validation/test split ratio of 48/32/20 as previous works [22, 35]. The results are shown in Table 5 and we can find that:
(1) When labels are sufficient, GraphACL and MUSE demonstrate impressive performance on both homophilous and heterophilous graphs. This can be attributed to their frameworks that adequately consider the heterophily on graphs.
(2) Self-Pro still outperforms traditional supervised models and self-supervised GNN models across most datesets, indicating that our framework can effectively transfer the pre-trained knowledge to downstream tasks once again.
(3) Most baselines perform even worse than simple MLP on heterophilous graphs, indicating the importance of feature semantic. This also confirms that our model can handle the heterophily on graphs through semantic prompt injection.
Table 6: Link prediction results on homophilous graphs. The best and second best performance for each dataset are marked with bold and underline.
@@ 5.4 Ablation Study (RQ3)
We still follow the above setting of few-shot node classification task. In Fig. 4 we can see the results:
(1) The degradation without prompt template mainly stems from the negative transfer problem, especially when the labeled nodes are limited.
(2) The decline without tuning indicates that the projector plays a crucial role on the knowledge transfer between pre-training and downstream tasks.
(3) Prompt injection improves the performance, indicating that providing extra information based on graph itself is beneficial for tasks with limited labels and the projector can effectively integrate the prompted information to enhance the performance in downstream tasks.
@@ 5.5 Parameter Analysis (RQ4)
We end this section with a sensitivity analysis on hyper-parameters in Self-Pro, i.e., the weight of prompt injection $\mu$. Meanwhile, we demonstrate the results using self weight $w$ mentioned in Section 4.4. Specificlly, we conduct experiments on node classification by varying $\mu$ from 0 to 1 to compare with the self weight strategy. Fig. 5 illustrates the accuracy under different $\mu$ on Cora, Citeseer and Texas. We see that self weight strategy and smaller injection weights can assist in node classification on homophilous graph, while larger injection ratios are beneficial for heterophilous graphs where semantic information is more crucial.
@@ 6 Conclusion
In this paper, we proposed Self-Pro, a framework utilizing self-prompting and tuning mechanism for graphs. The framework incorporated asymmetric graph contrastive learning to capture complex patterns for graphs with homophily and heterophily and aligned the pretext and downstream tasks into the same template. Then we reused the projector in the pre-training phase as the downstream adapter for efficient tuning and injected prompts from semantic and structure perspectives based on the information inherented in the graph itself. Finally, our experimental results demonstrated the superiority of Self-Pro.
7 Acknowledgement
This work is supported by National Natural Science Foundation of China No. 62202172 and Shanghai Science and Technology Committee General Program No. 22ZR1419900.